An object with a density of 941.0 kg/m3 and a mass of 895.0 kg is thrown into the ocean. Find the volume that sticks out of the water. (use ?seawater = 1024 kg/m3)
Answer #1

Mass of object m = 895 kg

density of object d = 941 kg/m3

density d = Mass / Volume = m/V

dencity of seawater dw = 1024 kg/m3

Thus, volume of object:

V = m/d = 895/941 = 0.951 m3

Now according to Archimedes principle, when an object is thrown in water, it displaces the amount of water equal to its mass:

Thus the amount or volume of water having mass same as that of object:

Vw = m/dw = 895 / 1024 = 0.874 m3

This is the amount of water displaced by the object and volume taken up by it when it is thrown in water.

Now only 0.874 m3 of its volume is in water.

Thus volume sticking out is:

V' = V - Vw= 0.951 - 0.874 = 0.77 m3